# Anti-Degeneracy Scheme for Lidar SLAM based on Particle Filter in Geometry Feature-Less Environments

### v1

@@ Abstract
Simultaneous localization and mapping (SLAM) based on particle filtering has been extensively employed in indoor scenarios due to its high efficiency. However, in geometry feature-less scenes, the accuracy is severely reduced due to lack of constraints. In this article, we propose an anti-degeneracy system based on deep learning. Firstly, we design a scale-invariant linear mapping to convert coordinates in continuous space into discrete indexes, in which a data augmentation method based on Gaussian model is proposed to ensure the model performance by effectively mitigating the impact of changes in the number of particles on the feature distribution. Secondly, we develop a degeneracy detection model using residual neural networks (ResNet) and transformer which is able to identify degeneracy by scrutinizing the distribution of the particle population. Thirdly, an adaptive anti-degeneracy strategy is designed, which first performs fusion and perturbation on the resample process to provide rich and accurate initial values for the pose optimization, and use a hierarchical pose optimization combining coarse and fine matching, which is able to adaptively adjust the optimization frequency and the sensor trustworthiness according to the degree of degeneracy, in order to enhance the ability of searching the global optimal pose. Finally, we demonstrate the optimality of the model, as well as the improvement of the image matrix method and GPU on the computation time through ablation experiments, and verify the performance of the anti-degeneracy system in different scenarios through simulation experiments and real experiments.
This work has been submitted to IEEE for publication. Copyright may be transferred without notice, after which this version may no longer be available.
@@
Figure 1: The schematic shows the state of two particle swarms before and after anti-degeneracy improvement.
@@ I Introduction
Slam is the core of realizing navigation for robots, which enables devices to perform real-time localization and map construction in unknown environments. By utilizing SLAM technology, devices such as service robots and automated guided vehicles are able to accurately locate themselves based on the environment map, providing the necessary preconditions for path planning and obstacle avoidance.
However, in geometry feature-less scenarios, such as long straight corridors, underground pipelines and spacious halls, the accuracy of SLAM is relatively low. This is due to the sparseness of the environmental features causes SLAM to lack sufficient constraints for accurate self-localization.
In monotonous color walls or scenes with repetitive textures, camera-based SLAM systems have difficulty extracting enough feature points for tracking, resulting in degeneracy of positioning accuracy and map quality. In addition, in low-light or strongly reflective conditions, the quality of the image captured by the camera can also be significantly degraded, affecting feature extraction and depth estimation [1, 2, 3, 4, 5, 6]. Lidar-based SLAM is difficult to perform effective feature matching in degraded environments due to insufficient captured point cloud features. Especially in long straight corridors, the lack of orthogonal orientation constraints makes it difficult for the robot to perform accurate localization [7, 8, 9, 10].
Particle swarm distribution is the most intuitive representation of SLAM convergence state, so it is different from other degeneracy detection work based on SLAM process eigenvalues, in this article, we propose a detection method based on particle distribution, and design a self-tuning optimization strategy based on the degree of degeneracy. Compared to other robust SLAM methods with fixed optimization strategies like [11], our method can adaptively adjust the optimization strategy based on the level of degeneracy. This flexibility allows the system to better adapt to different environmental conditions. First, it uses linear mapping and Gaussian augmentation to obtain training samples of degeneracy particles. Through supervised learning, a neural network is endowed with degeneracy detection capabilities. Subsequently, the framework employs the model to analyze particle distribution to detect the degree of degeneracy. Finally, we use an optimization framework based on the degree of degeneracy, which adaptively adjusts strategy to enhance the global optimality of pose estimation and change the contributions of different sensors. The main contributions of this article are summarized as follows:
@@ II-B Anti-Degeneracy Strategies
Anti-degeneracy strategy is the key element to improve SLAM accuracy in degraded environments. Li et al. [17] dealt with the degeneracy problem through geometric and intensity-based feature extraction by designing two multi-weight functions to fully extract the features of planar and edge points. Zhang et al. [18] presented an innovative tightly coupled lidar-vision-inertial odometry, which enables accurate state estimation in degraded environments through sensor fusion strategies. Zhang et al. [19] utilized ARTag as a visual marker to assist in localization, which employs a bitmap optimization method to reduce the error. Lee et al. [12] proposed a switching-based SLAM that achieves high accuracy by switching from lidar to visual odometry upon detection of lidar degeneracy. However, these methods have requirements on the geometry of environment and the strength of features. Meanwhile, the multi-sensor fusion strategy increases the deployment cost, the parameters need to be adjusted for adaptation, which leads to the fact that these methods do not have good robustness.
@@ III System overview
The framework diagram of the anti-degeneracy system presented in this article is shown in Fig. 2. Particle filter-based SLAM represents a posteriori probability distribution of robot positions and maps by maintaining a set of random particles [20, 21]. Particles are first preliminarily updated using the odometry data from the IMU and motor encoder through the motion model, as shown in (1). Subsequently, the observation model is employed to update the particles based on the lidar data to determine the optimal pose through pose estimation. The particle weights are then calculated according to (2), reflecting the uncertainty of the pose estimation. Finally, resampling and map updating operations are carried out.
@@
Degeneracy is typically manifested as the uncertainty of the pose, which is reflected by the weights and distribution of the particles. Dense particle swarms can better capture the nonlinear characteristics of the state space and reduce the effect of random errors. In feature-rich environments, the particle swarms tend to be aggregated, which represents high localization accuracy, conversely, in degraded environments, the particle swarms tend to be dispersed, which represents low accuracy, as shown in Fig. 1. To this end, we transform the degeneracy detection into a classification task for a neural network, analyzing the degree of degeneracy through the dispersion of particles.
We utilize robot operating system (ROS) to transmit particle coordinates to the degeneracy detection node, where they are converted into images via linear mapping and processed using data augmentation based on a Gaussian model. After the model predicts the degree of degeneracy, the anti-degeneracy node receives the prediction feedback via ROS and adaptively adjusts the optimization strategy according to the degeneracy level to enhance the search for the optimal pose and dynamically adjust the weights of different sensors.
Figure 2: Overall flow chart of the system.
Figure 3: The architecture of degeneracy detection model.
@@ IV-A1 Linear Mapping with Scale Invariance
The quality of the particle image dataset is crucial for the model training. The uncertainty in the coordinate size of the particles leads to them being stretched or shrunk in various dimensions, making the relative position and distribution of the particles distorted. In order to solve this problem, we design a mapping method from continuous space to discrete index for particle swarms of various sizes that ensures they maintain original distribution and result in scale-invariant images.
We first define the canvas size of the image as 10m in reality, which defines the accuracy threshold for localization and the outliers are considered invalid particles. Subsequently, the image boundary are modified according to (3). The procedure is to first modify the extreme values of the particle coordinates in the x and y dimensions, ensuring that the coordinates in each dimension have the same discrete mapping metric respectively.
@@
Then we discretize the continuous spatial coordinates of the particle swarm into pixel indices, translating physical locations to image space pixels. By linearly mapping the coordinate extremes changes in (3) to all particles, the scale invariance can be matained across both the physical and image spaces, as detailed in (4), where $i$ represents the discrete index corresponding to the particle coordinates in the image, which represents the pixel position.
@@ IV-A2 Data Augmentation based on Gaussian Model
Particle distributions can be affected by changes in number, which in turn affects model performance. For example, particle swarm in a clustered state may become sparse due to a decrease in number. In addition, too many particles can lead to overlapping problem, which can also change the original distribution.
In order to address this problem, we adopt the Gaussian model to simulate the influence of each particle on the surrounding like mesh, where the position of the particle serves as the center of the Gaussian distribution, and the influence of the particle decreases exponentially with distance. The Gaussian function is defined as:
@@
where $(x,y)\in\mathbb{Z}^{2}$ and $(x_{p},y_{p})\in\mathbb{Z}^{2}$ represent the grid positions influenced by the particles and of the particles themselves, respectively; and $\sigma$ is the standard deviation, we adjust it to make the influence range of the Gaussian be a 5*5 grid, as we found that the augmentation effect of 5*5 grid is good for the commonly used number of particles.
For each grid, we compute the Gaussian contributions of other particles to it and then accumulate these contributions to update the occupancy of that grid:
@@
where $o$ stands for the occupancy, we sum Gaussian influence within a 5*5 range around each grid. If the value exceeds a threshold, it is considered occupied. We prefer a fixed low threshold, as more occupied grids would not affect the result.
@@ IV-A3 Degeneracy Detection Model
As the number of particles increases, the model needs to deepen its architecture to capture complex features. Compared to other classical models, the design based on ResNet can more flexibly expand its depth. After introducing transformer, the accuracy is increased by about 5%.
The overall architecture of the model is shown in Fig. 3. After obtaining the image dataset of the particle swarm, we first grayscale the image to preserve the contour and shape of the particle swarm and remove redundant information. Subsequently, the grayscaled data first passes through a convolutional layer with a convolutional kernel size of 7*7 and a step size of 2, which facilitates the extraction of abstract features by the model. Then it passes through a batch normalization (BN) layer, which solves the problem of bias of internal covariates in the network by regularization method. This process involves calculating the mean and variance of the input feature x, as shown in (7). Then the features are normalized using (8), with a $\epsilon$ added to the denominator to prevent division by zero. Finally introduce the learnable parameters $\gamma$ and $\beta$, and perform the scaling and panning according to (9). This network layer can speed up the convergence of the model and alleviate the problem of exploding or vanishing gradients. The next layer is the activation function layer, where we employ the ReLU function to enhance the nonlinear representation of the model. This is followed by a maximum pooling layer for downsampling to further reduce the computational effort.
$\begin{cases}\mu=\frac{1}{N}\sum_{i=1}^{N}x_{i}\\ \\ \sigma^{2}=\frac{1}{N}\sum_{i=1}^{N}(x_{i}-\mu)^{2}\\ \end{cases}$ (7)
@@
Next, based on the ResNet architecture, we design the backbone of the network, which consists of a total of 4 layers, and each layer gradually extracts the features of the image through a different number of BasicBlocks, and the abstraction level of the particle distribution features is gradually increased from the shallow layer to the deeper layer. The BasicBlock is a residual block structure, which contains 3*3 convolutional layers , BN, and ReLU functions, the convolution layer can reduce the spatial size of the feature map $f\in\mathbb{R}^{C\times H\times W}$, making the feature map more compact and highlighting the important features. BasicBlock also maintains the gradient flow by constant mapping to make the features skip certain network layers for propagation. The use of BN and ReLU activation function in it makes the network training more stable, accelerates the convergence, improves the generalization ability, and, finally, generates feature maps enriched with high-level semantic information and ensures consistent feature dimensions when the residuals are connected by downsample layer.
We take the transformer as the neck part of the network, and firstly, the feature map with the size of 256*28*28 output from the backbone is spread, and then the features are adjusted to match the dimensions of the input from the transformer through a linear layer. The adjusted features, which are input to the transformer encoder, are further enhanced through the self-attention mechanism and feed-forward network processing sequences to enable the model to capture long range dependencies and improve the representation of the features.
Finally, in the head portion of the network, we use a fully connected layer as a classification header, which is capable of converting high-dimensional features extracted from complex network structures into class-specific predictions, and the output is a vector consisting of two values ranging from 0 to 1 to describe the degree of degradation.
@@ IV-B1 Global Position Optimization Based on the Degree of Degeneracy
The position is first initially adjusted using the odometry model, and then the scan matching of observation model is used for further optimization. We divide the lidar-based position optimization into a combination of coarse and fine matching. In the coarse matching phase, the robot will do displacements in eight directions with a step size of $\delta_{l}\in\mathbb{R}$ and rotations with an angle of $\delta_{a}\in\mathbb{S}^{1}$, and after each action has been performed, the current matching score $S\in\mathbb{R}^{+}$ will be calculated, the one with the highest score will be selected as the best position and updated, whereas in the fine matching phase, the scales of $\delta_{l}/2$ and $\delta_{a}/2$ will be used to perform the same process. The number of executions of the above matching combinations will adaptively change with the degeneracy level to ensure the accuracy of the search space. In addition, we design an factor $\varphi$ that includes degeneracy level $c$ to control the dependence on different sensors by changing the matching score, which tends to trust the position given by the odometry when the lidar degeneracy is severe, as shown in (10).

### v2

@@ Abstract
Simultaneous localization and mapping (SLAM) based on particle filtering has been extensively employed in indoor scenarios due to its high efficiency. However, in geometry feature-less scenes, the accuracy is severely reduced due to lack of constraints. In this article, we propose an anti-degeneracy system based on deep learning. Firstly, we design a scale-invariant linear mapping to convert coordinates in continuous space into discrete indexes, in which a data augmentation method based on Gaussian model is proposed to ensure the model performance by effectively mitigating the impact of changes in the number of particles on the feature distribution. Secondly, we develop a degeneracy detection model (DD-Model) using residual neural networks (ResNet) and transformer which is able to identify degeneracy by scrutinizing the distribution of the particle population. Thirdly, an adaptive anti-degeneracy strategy is designed, which first performs fusion and perturbation on the resample process to provide rich and accurate initial values for the pose optimization, and use a hierarchical pose optimization combining coarse and fine matching, which is able to adaptively adjust the optimization frequency and the sensor trustworthiness according to the degree of degeneracy, in order to enhance the ability of searching the global optimal pose. Finally, we demonstrate the optimality of model, as well as the improvement of image matrix method and GPU on the inference time through ablation experiments. The effectiveness of our method is verified by comparing with SOTA methods.
@@
Figure 1: Anti-degeneracy strategy restores observation distribution to an approximate Gaussian distribution, improving the search for the optimal pose.
@@ I Introduction
Slam is the core of realizing navigation for robots, which enables devices to perform real-time localization and map construction in unknown environments. 2D-SLAM is widely used in the localization task of service robots and vacuum cleaners because of its low cost and high reliability [1, 2], and its output two-dimensional grid map is suitable for path planning and obstacle avoidance.
However, in geometry feature-less scenarios, such as long straight corridors, underground pipelines and spacious halls, the accuracy of SLAM is relatively low. This is due to the sparseness of the environmental features causes SLAM to lack sufficient constraints for accurate self-localization. In monotonous color walls or scenes with repetitive textures, camera-based SLAM systems have difficulty extracting enough feature points for tracking, resulting in degeneracy of positioning accuracy and map quality. In addition, in low-light or strongly reflective conditions, the quality of the image captured by the camera can also be significantly degenerate, affecting feature extraction and depth estimation [3, 4, 5, 6, 7, 8, 9]. Lidar-based SLAM is difficult to perform effective feature matching in degenerate environments due to insufficient captured point cloud features. Especially in long straight corridors, the lack of orthogonal orientation constraints makes it difficult for the robot to perform accurate localization [10, 11, 12, 13].
Shi et al. [14] proposed a degeneracy detection method for 2D-SLAM, he achieved detection by extracting normal features from Lidar point clouds and using principal component analysis for feature aggregation and analysis. However, this method needs to adjust parameters in different scenarios. The challenge of degeneracy detection is to improve the generalization ability in different scenarios and get rid of the dependence on environmental conditions and preassumptions.
Particle swarm distribution is the most intuitive representation of SLAM convergence state, as shown in Fig. 1, this distribution pattern can be captured by the neural network to achieve highly generalized degeneracy detection. In this article, we propose a detection method for particle-filtering based 2D-SLAM, and design a self-tuning optimization strategy based on the degree of degeneracy. Compared to other robust SLAM methods with fixed optimization strategies like [15], our method can adaptively adjust the optimization strategy based on the level of degeneracy. This flexibility allows the system to better adapt to different environmental conditions. First, it uses linear mapping and Gaussian augmentation to obtain training samples of degeneracy particles. Through supervised learning, a neural network is endowed with degeneracy detection capabilities. Subsequently, the framework employs the model to detect degeneracy by analyzing particle distribution. Finally, we use an optimization framework based on the degree of degeneracy, which adaptively adjusts strategy to enhance the global optimality of pose estimation and change the contributions of different sensors. The main contributions of this article are summarized as follows:
@@ II-B Anti-Degeneracy Strategies
Anti-degeneracy strategy is the key element to improve SLAM accuracy in degenerate environments. Li et al. [21] dealt with the degeneracy problem through geometric and intensity-based feature extraction by designing two multi-weight functions to fully extract the features of planar and edge points. Zhang et al. [22] presented an innovative tightly coupled lidar-vision-inertial odometry, which enables accurate state estimation in degenerate environments through sensor fusion strategies. Zhang et al. [23] utilized ARTag as a visual marker to assist in localization, which employs a bitmap optimization method to reduce the error. Lee et al. [16] proposed a switching-based SLAM that achieves high accuracy by switching from lidar to visual odometry upon detection of lidar degeneracy. However, these methods have requirements on the geometry of environment and the strength of features. Meanwhile, the multi-sensor fusion strategy increases the deployment cost, the parameters need to be adjusted for adaptation, which leads to the fact that these methods do not have good robustness.
@@ III System overview
The framework diagram of the anti-degeneracy system presented in this article is shown in Fig. 2. Particle filter-based SLAM represents a posteriori probability distribution of robot positions and maps by maintaining a set of random particles [1]. Particles are first preliminarily updated using the odometry data from the IMU and motor encoder through the motion model, as shown in (1). Subsequently, the observation model is employed to update the particles based on the lidar data to determine the optimal pose through scanmatch of pose estimation. If the environment features are rich enough for scanmatch to succeed, the particle weights which reflect the uncertainty of pose estimation are then calculated according to (2). Finally, resampling and map updating operations are carried out.
@@
As shown in Fig. 1, in non-degenerate case, SLAM has sufficient environmental constraints. The second order taylor expansion is performed on observation likelihood $p(z_{t}|x_{t})$ in (3), where $I_{o}$ is observation fisher matrix, $x^{*}$ is groudtruth of pose, $z_{t}$ and $x_{t}$ are lidar beam and pose at time $t$ respectively:
$\log p(z_{t}|x_{t})\approx\log p(z_{t}|x_{t}^{*})-\frac{1}{2}(x_{t}-x_{t}^{*})^{T}I_{o}(x_{t}-x_{t}^{*})$ (3)
At this moment, particle distribution radius is constrained by the Mahalanobis distance shown in (4), where $\Sigma=I_{o}^{-1}$ is posterior covariance matrix , $\chi_{d,1-\alpha}^{2}$ is the critical value of the Chi-square distribution when the degree of freedom is $d$ and the significance level is $a$:
$(x_{t}-x_{t}^{*})^{T}\Sigma^{-1}(x_{t}-x_{t}^{*})\leq\chi_{d,1-\alpha}^{2}$ (4)
In this case, all dimensions of covariance are small, and particles cluster around $x^{*}$. Each particle has a high fitting probability for groundtruth. The dense distribution indicates that the sampling space of pose optimization is relatively certain, which can capture the nonlinear characteristics of the state space and improve the localization accuracy.
In degenerate case, the robot lacks environmental constraints in the travel direction. Particles obtain similar likelihood scores via scanmatch in this direction, causing scanmatch to fail and failing to constrain particle updates. The likelihood distribution $p(z_{t}|x_{t})$ diverges in the travel direction and degenerates from gaussian distribution to flat distribution shown in (5), where $\epsilon$ is the eigenvalue of $I_{o}$ in the direction of travel:
$p(z_{t}|x_{t})\propto exp(-\frac{1}{2}\epsilon(x-x^{*})^{2})$ (5)
Then particles cover a wider state space to avoid falling into local optimality and alleviate localization stagnation to some extent. Conversely, in the direction perpendicular to travel, the homogenization of lidar ranging information is severe. Although the constraint information is inaccurate, the severity of the degeneracy problem in the travel direction dominates the divergence direction of particles. This results in compression of the distribution in the perpendicular direction, reducing the fitting probability to the ground truth (GT). The error $\Delta x$ in the direction of travel is projected into the vertical direction through the turning angle $\theta$, causing the map to skew:
$\Delta y\approx\Delta x\cdot sin\theta$ (6)
We transform the degeneracy detection into a classification task for a neural network, analyzing the degeneracy through the dispersion of particles. We utilize robot operating system (ROS) to transmit particle coordinates to the DD-Model, where they are converted into images via linear mapping and processed using data augmentation based on gaussian, and then the model predicts the degree of degeneracy. If the robot is in a degenerate environment, then scanmatch fails, as shown in Fig. 2, the anti-degeneracy node will improve the pose optimization by hierarchical anti-degeneracy strategies according to the degeneracy level, the likelihood distribution is approximately converted to gaussian form to increase the search capability for the globally optimal pose, and the contribution of different sensors is adjusted by influencing the likelihood score.
Figure 2: Overall flow chart. Anti-degeneracy strategy will be triggered when scanmatch fails due to degeneracy, increasing the hierarchical pose optimization count by one for every 0.25 increase in degeneracy.
Figure 3: The architecture of DD-Model.
@@ IV-A1 Linear Mapping with Scale Invariance
The quality of the particle image dataset is crucial for model training. Uncertainty in particle coordinates can distort their relative positions and distribution. We design a mapping method from continuous space to discrete indices for particle swarms of various sizes, ensuring they maintain their original distribution and produce scale-invariant images.
We first define the canvas size of image as 10m in reality, which defines the accuracy threshold for localization and the outliers are considered invalid particles. The increasing number of particles that exceed the boundary leads to a progressively sparser distribution, thereby augmenting the probability that the model will classify the situation as degenerate, consistent with the anticipated outcome. Subsequently, the image boundary are modified according to (7). The procedure is to first modify the extreme values of the particle coordinates in the x and y dimensions, ensuring that the coordinates in each dimension have the same discrete mapping metric respectively.
@@
Then we discretize the continuous spatial coordinates of particles into pixel indices, translating physical locations to image space pixels. By linearly mapping the coordinate extremes changes in (7) to all particles, the scale invariance can be matained across both the physical and image spaces, as detailed in (8), where $i$ is discrete index corresponding to particle coordinates, indicating the pixel position in image.
@@ IV-A2 Data Augmentation based on Gaussian Model
Particle distributions can be influenced by changes in their number, impacting model performance. For instance, a clustered particle swarm may become sparse if the number decreases. Conversely, too many particles can cause overlap, altering the original distribution. We adopt the Gaussian model in (9) to simulate the influence of each particle on the surrounding grid, where the position of the particle serves as the center of the Gaussian distribution, and the influence of the particle decreases exponentially with distance.
@@
where $(x,y)\in\mathbb{Z}^{2}$ and $(x_{p},y_{p})\in\mathbb{Z}^{2}$ represent the grid positions influenced by the particles and of the particles themselves, respectively, and $\sigma$ is the standard deviation. We adjust it to make the Gaussian influence range a 5*5 grid, as this size optimizes augmentation effects for the typical number of particles and other size range will reduce the effect.
For each grid, we accumulate the contribution of particles to update the occupancy $o$ of this grid:
@@
If $o$ exceeds a preset threshold, the grid is marked as occupied. Since it should mark more grids as occupied when overlap is severe, we prefer to set a lower threshold. Notably, when the number of particles is too low, the method will only mark grids in dense particle areas as occupied, while sparse areas remain unchanged, achieving the correct augmentation effect.
@@ IV-A3 Degeneracy Detection Model
The overall architecture of the model is shown in Fig. 3. We first grayscale the image to preserve the contour and shape of the particle swarm and remove redundant information. Subsequently, the data passes through a convolutional layer with a 7*7 kernel, which facilitates the extraction of features. Then it passes through a batch normalization (BN) layer with ReLU, which solves bias problem of internal covariates by regularization and enhance the ability of nonlinear representation. Specifically, features are normalized using (11), with $\epsilon$ added to the denominator to avoid division by zero. Then, scaling and shifting are performed using learnable parameters $\gamma$ and $\beta$ as (12). Then a maximum pooling layer follows for downsampling, further reducing the computational effort.
@@
As the scale and complexity of particle swarm increase, model accuracy tends to decline, and deepening the network architecture to improve accuracy often influence convergence effect during training. ResNet effectively addresses these gradient-related problems with residual learning and skip connections while reducing parameter count and inference time. We design backbone based on ResNet, which consists of four composite layers. Each layer extracts features using a varying number of BasicBlocks. The BasicBlock is a residual block structure, which contains 3*3 convolutional layers and BN layers with ReLU. The convolutional layer reduce the spatial size of feature map $f\in\mathbb{R}^{C\times H\times W}$ and highlight important features, increasing the abstraction level of features from shallow to deeper layer. BasicBlock also maintains gradient flow by constant mapping to make the features skip certain layers for propagation, accelerating convergence and improving generalization ability, and, finally, generates feature maps that are enrich with high-level semantic information and match the dimensions of downsampling layer.
We take transformer as the neck. First, 256*28*28 feature map output from backbone is spread, and then features are adjusted to match the dimension of transformer encoder through linear layers. The adjusted features are further enhanced through self-attention mechanism and feed-forward network sequences, to enable model to capture long range dependencies and improve the representation of features. After introducing transformer, the accuracy is increased by about 5%.
Finally, we use a fully connected layer as the classification header, which is capable of converting high-dimensional features into class-specific predictions, and the output is a vector consisting of two values ranging from 0 to 1 to describe the degree of degeneracy.
@@ IV-B1 Global Position Optimization Based on the Degree of Degeneracy
The position is first initially adjusted using the odometry model, and then the scan matching of observation model is used for further optimization. We divide the lidar-based position optimization into a combination of coarse and fine matching. In the coarse matching phase, the robot will do displacements in eight directions with a step size of $\delta_{l}\in\mathbb{R}$ and rotations with an angle of $\delta_{a}\in\mathbb{S}^{1}$, and after each action has been performed, the current matching score $S\in\mathbb{R}^{+}$ will be calculated, the one with the highest score will be selected as the best position and updated, whereas in the fine matching phase, the scales of $\delta_{l}/2$ and $\delta_{a}/2$ will be used to perform the same process. The number of executions of the above matching combinations will adaptively change with the degeneracy level to ensure the accuracy of the search space. The score of scanmatch directly determines the credibility of the current pose. We scale the score of lidar through an factor $\varphi$ that includes degeneracy level $c$ in (13), so that the score of lidar is lower when the degeneracy is serious. At this time, the score of odometry pose is higher and closer to the real pose, so the odometry pose will be selected as the optimal pose. All of the above methods essentially achieve degeneracy optimization by searching for poses with higher scores, avoiding the explicit dependence on the direction of degeneracy.